OpenStudy (anonymous):

A sphere has a diameter of 4 meters. What is the approximate volume of the sphere?

OpenStudy (unklerhaukus):

did you know that the formula for the volume of a sphere is \[V_{\text{sphere}}=\frac 43 \pi r^3\] you'll also have to use \(d=2r\) or \(r=\frac d2\)